OpenStudy (anonymous):

\[\int_0^\infty\frac{x^8-4x^6+9x^4-5x^2+1}{x^{12}-10x^{10}+37x^8-42x^6+26x^4-8x^2+1}dx\]

OpenStudy (agreene):

\[\int_0^\infty \frac{x^7-4x^5+9x^3-5x}{x^{11}-10x^9+37x^7-42x^5+26x^3-8x+1}\] Thats about all I can see to do at this point. It definitely seems doable, just a bit of a headache... wolfram says it is ultimately \[= \frac{\pi}{2}\]

OpenStudy (anonymous):

I hope this doesn't show up in any test :-(

OpenStudy (agreene):

well, more than likely, you are going to need to split it into all it's fractional parts and then start doing substitution tricks to get all the individual integrals to work out.

OpenStudy (agreene):

Only problem is, when that method doesnt work... you've wasted a good bit of your life, lol.

OpenStudy (anonymous):

I'm glad I did well on that integration quiz

OpenStudy (anonymous):

I got away with an A

OpenStudy (agreene):

Nice.